# scenarioparampadding=8pt, margin-top=3pt, margin-bottom=3pt, border-style=none, background-color=quotebackground

### v1

@@ 6.4. MHPs’ Perspectives on MindfulDiary for Clinical Mental Health Settings
6.4.1. Expert Usage of MindfulDiary
In this section, we described how MHPs utilized the expert interface and what expectations and limitations they perceived. For this purpose, we interviewed five psychiatrists who were in charge of outpatient treatment at University Hospital.
All of the psychiatrists emphasized the critical value of an expert interface based on information recorded in the daily lives of patients. Specifically, E3 highlighted the program’s value in that MindfulDiary consistently aids in recording daily entries, allowing them to utilize more detailed patient data during outpatient visits. “Patients, with the support of AI, can logically continue their narratives, ensuring more dialogue than a typical (paper-based) diary. This definitely aids me in my consultations.” (E3). In this section, we present some examples of how MindfulDiary was integrated into clinical settings.
Integrating MindfulDiary into Clinical Workflow. First, we investigate how psychiatrists use MindfulDiary in their workflow. We have not provided guidance on how to use the clinician dashboard. However, we informed psychiatrists that the LLM-based analysis may not be accurate, and therefore, it is necessary to check the interaction log-data. The psychiatrists reviewed MindfulDiary data from their patients every morning during the deployment study, which they typically spent reviewing patient chart data they were going to see that day. Depending on the severity and the focal concerns of the patient, psychiatrists spent about 5-10 minutes per patient reviewing the MindfulDiary data. After observing trends primarily through PHQ-9 in the clinician dashboard first, psychiatrists read summaries about overall events and documented emotions. If there were spikes or drops in the PHQ-9 or events/emotions that needed further verification, they checked the actual interaction log.
@@ 6.4.1. Expert Usage of MindfulDiary
6.4.2. Perceived Opportunities and Limitations of MindfulDiary
While mental health MHPs generally appraised the utility of the MindfulDiary positively, they also perceived potential limitations, sharing considerations when integrating MindfulDiary into clinical settings.
@@ 7.1. Guiding patient journaling through conversations offering diverse perspectives
Our study highlighted the potential of MindfulDiary in clinical settings, mainly where adherence to interventions is important (Musiat et al., 2022). Core symptoms of depression, such as loss of energy, difficulty in carrying out mental processes, and feelings of apathy, often contribute to lower adherence to a professional’s advice or intervention (Kennedy, 2022). Clinicians who participated in our FGI also highlighted these challenges in motivating patients to utilize the diary writing program. Our findings demonstrated that MindfulDiary helped mitigate these challenges by transforming the conventional journaling process into engaging conversations. Using MindfulDiary, users were able to engage in conversations with the system by answering prompts and questions, which made them feel the journaling process was more accessible and intriguing. This active participation ensures that the users are not overwhelmed by the task and are guided in documenting their feelings and experiences more richly.
@@
7.2. Chatbot as a Mediator for Fostering Patient Understanding
Studies have suggested that sharing the data captured via chatbots with others, such as health professionals and family members, could further serve as an effective mediator that helps convey more truthful information (Lukoff et al., 2018; Lee et al., 2020a). For instance, patients consistently displayed deep self-disclosure through chatbots, whether or not they intended to share their inputs with health professionals (Lee et al., 2020a). Aligned with prior work on PGHD (Cohen et al., 2016; Nundy et al., 2014), MHPs in our study also perceived that MindfulDiary has shed light on patients’ daily events, emotions, and thoughts that might have been difficult to gain through regular clinical visits. This data offered MHPs valuable insights into the patient’s experiences and context. However, our study findings point to the need to carefully consider. In our study, we observed that the act of sharing journal content with MHPs through MindfulDiary led some patients to exaggerate their conditions or needs. This case underscores that when integrating a system like MindfulDiary into a clinical context, not only the design of chatbots that lead to patient disclosure behavior but also the complex patient-provider dynamics in clinical settings play a significant influence. The growing prevalence of chatbots in mental health domains emphasizes the need for a holistic approach to their design and implementation. We highlight that engineers and MHPs need to collaborate closely, ensuring that these tools are not only technically sound but also tailored to meet the intricate dynamics of clinical settings (Thieme et al., 2022).
7.3. Considerations for Integrating LLM-driven Patient Program
In this section, we discuss the consideration for integrating LLM-driven patient programs into clinical mental health settings, drawing insights from the design and evaluation of MindfulDiary.
@@ 7.3. Considerations for Integrating LLM-driven Patient Program
Tailored LLM Evaluation for Clinical Mental Health Domains. The domain of mental health, which our study addresses, is characterized by the vulnerability of its target user group. The content discussed within this domain is often emotionally charged and sensitive. Therefore, prioritizing user safety becomes even more essential in this domain than in others. Considering the sensitivity of the domain, during our evaluation process, MHPs thoroughly tested the LLM’s output by trying out conversations on various sensitive topics in both implicit and explicit ways, drawing upon their clinical experiences. The contents the MHPs input were much more diverse and wide-ranging than what engineers could generate during the development. Additionally, MHPs showed concern that the hallucinations of the language model could reinforce or expand the delusions of patients with delusional disorders. We highlight that developing evidence-based tests or benchmark sets to anticipate the behavior of the language models in collaboration with MHPs is critical when leveraging LLMs for clinical mental health settings.
Role of MHPs in Integration. Considering the limitations of current LLMs (Korngiebel and Mooney, 2021), it is critical to involve MHPs when deploying LLM-driven systems for patients in mental health contexts. While planning the field deployment study of MindfulDiary, we identified specific roles that MHPs could play. In the pre-use phase, MHPs should determine the suitability of users and facilitate the onboarding process with patients. During the mid-use phase, they should closely monitor interactions with the LLM and be prepared to intervene in cases of crises or unexpected use scenarios. Furthermore, they can offer or adjust treatments based on long-term data periodically. Additionally, they should regularly re-evaluate the continued use of the system. While some of these tasks should carefully be designed not to burden MHPs too much, it is important that LLMs do not make autonomous decisions about patients (e.g., diagnosis, prescription, or crisis management) but instead operate under professional oversight.
@@ 7.4. Limitations and Future Work
Our recruitment method could impact the generalizability of our findings, as we recruited the patient participants for our field study from a single university hospital. Although we aimed to recruit patients with diverse types and levels of symptoms, our participants are not representative samples of psychiatric patients. They were young (mostly adolescents) and consulted by a fixed number of psychiatrists. While this work is just a first step toward designing an LLM-driven journaling app for psychiatric patients, further investigation is necessary with subjects from various backgrounds.
To implement our pipeline, we used OpenAI’s GPT API, which provided the most capable LLM at the time of our study and was accessible via commercial API. As GPT models are continually updated, later models may not yield the same conversational behavior. To generalize the performance of our conversational pipeline design, future work is needed to compare multiple versions of MindfulDiary with different underlying LLMs.
@@ Acknowledgements.
We thank our study participants for their time and efforts. We also thank Eunkyung Jo and Yubin Choi for providing feedback on the early draft of this paper. This work was supported as a research internship at NAVER AI Lab.

### v2

@@ 6.4. MHPs’ Perspectives on MindfulDiary for Clinical Mental Health Settings
In this section, we describe how MHPs utilized the clinician dashboard and the benefits and drawbacks of the system they reported, drawing on the debriefing interviews with the psychiatrists.
6.4.1. Utilization of MindfulDiary in Clinic
During the deployment study, psychiatrists reviewed the journal entries from their patients every morning when they reviewed the medical charts of patients whom they would meet on the day. Depending on the severity and the focal concerns of the patient, psychiatrists spent about 5 to 10 minutes per patient reviewing the MindfulDiary data. After checking trends primarily through PHQ-9 in the clinician dashboard, psychiatrists read summaries about events and documented emotions. If there were spikes or drops in the PHQ-9 or events/emotions, they checked the actual dialogues.
6.4.2. Percevied Benefits of MindfulDiary for Enhanced Patient Insight and Empathetic Engagement
All of the psychiatrists emphasized the critical value of an expert interface based on information recorded in the daily lives of patients. Specifically, E3 highlighted MindfulDiary’s value in that it consistently aids in recording daily entries, allowing them to utilize more detailed patient data during outpatient visits. “Patients, with the support of AI, can logically continue their narratives, ensuring more dialogue than a typical (paper-based) diary. This definitely aids me in my consultations.” (E3). In this section, we further report on how MindfulDiary has been helpful in the clinical practice of psychiatrists.
@@ 6.4.2. Percevied Benefits of MindfulDiary for Enhanced Patient Insight and Empathetic Engagement
6.4.3. Perceived Concerns about MindfulDiary
While MHPs generally appraised the utility of the MindfulDiary positively, they also raised concerns regarding the integration of MindfulDiary into clinical settings.
@@ 7.1. Guiding Patient Journaling through Conversations Offering Diverse Perspectives
Our study highlighted the potential of MindfulDiary in clinical settings, mainly where adherence to interventions is important (Musiat et al., 2022). Core symptoms of depression, such as loss of energy, difficulty in carrying out mental processes, and feelings of apathy, often contribute to lower adherence to a professional’s advice or intervention (Kennedy, 2022). Clinicians who participated in our FGI also highlighted these challenges in motivating patients to utilize the diary writing app. Our findings demonstrated that MindfulDiary helped mitigate these challenges by transforming the conventional journaling process into engaging conversations. Using MindfulDiary, users were able to engage in conversations with the system by answering prompts and questions, which made them feel the journaling process was more accessible and intriguing. This active participation ensures that the users are not overwhelmed by the task and are guided in documenting their feelings and experiences more richly.
@@
7.2. MindfulDiary as a Facilitator for Fostering Patient-Provider Communication
Studies have suggested that sharing the data captured via chatbots with others, such as health professionals and family members, could further serve as an effective mediator that helps convey more truthful information (Lukoff et al., 2018; Lee et al., 2020a). For instance, patients consistently displayed deep self-disclosure through chatbots, whether or not they intended to share their inputs with health professionals (Lee et al., 2020a). Aligned with prior work on PGHD (Cohen et al., 2016; Nundy et al., 2014), MHPs in our study also perceived that MindfulDiary has shed light on patients’ daily events, emotions, and thoughts that might have been difficult to gain through regular clinical visits. This data offered MHPs valuable insights into the patient’s experiences and context.
Building on these findings, we could expand the potential presented by MindfulDiary in patient-provider communication. In the field of personal health informatics, existing research highlights the role of technology, such as photo journaling, in managing conditions like Irritable Bowel Syndrome. This tool not only empowers patients to record their daily experiences more effectively but also fosters enhanced collaboration between patients and healthcare providers (Chung, 2017; Schroeder et al., 2017). Such tools serve as vital artifacts in negotiating the boundaries of patient-provider interactions (i.e., boundary negotiating artifacts) (Chung et al., 2016).
This work adds a new dimension to this discussion by showing how LLM-assisted journaling lowers barriers to generating health data in daily life and fosters patient understanding. Specifically, we found that through this system, patients and providers can collaboratively reflect on mental health conditions. In the context of the stage-based model of personal informatics, the patient module in our MindfulDiary helps patients reduce the burden of collecting daily data and supports deeper recording. The expert module’s dashboard allows for the combined and transformed processes of diary data, survey data, and quantitative engagement data, supporting MHPs’ integration and reflection. Collaborative data generation and utilization with patients can enable care that reflects the patient’s values and the characteristics of their daily life. These insights serve as a basis for patient-provider collaboration.
However, our study findings underscore the importance of careful consideration in the clinical integration of systems like MindfulDiary. While we did not observe patients exaggerating their conditions or needs, this potential issue was raised as a concern by MHPs. They expressed apprehension about the possibility that sharing journal content with MHPs through MindfulDiary might lead some patients to exaggerate their conditions or needs. This concern highlights the need to consider not only the design of chatbots that facilitate patient disclosure behavior but also the complex dynamics between patients and providers in clinical settings. It is crucial to address these dynamics to ensure the effective and safe use of such technologies in mental health care. The growing prevalence of chatbots in mental health domains emphasizes the need for a holistic approach to their design and implementation. We highlight that engineers and MHPs need to collaborate closely, ensuring that these tools are not only technically sound but also tailored to meet the intricate dynamics of clinical settings (Thieme et al., 2022).
7.3. Considerations for Integrating LLMs into Clinical Settings
In this section, we discuss the consideration for integrating LLMs into clinical mental health settings, drawing insights from the design and evaluation of MindfulDiary.
@@ 7.3. Considerations for Integrating LLMs into Clinical Settings
Tailored LLM Evaluation for Clinical Mental Health Domains. The domain of mental health, which our study addresses, is characterized by the vulnerability of its target user group. The content discussed within this domain is often emotionally charged and sensitive. Therefore, prioritizing user safety becomes even more essential in this domain than in others. Considering the sensitivity of the domain, during our evaluation process, MHPs thoroughly tested the LLM’s output by trying out conversations on various sensitive topics in both implicit and explicit ways, drawing upon their clinical experiences. The contents the MHPs input were much more diverse and wide-ranging than what engineers could generate during the development. Additionally, MHPs showed concern that the hallucinations of the LLM could reinforce or expand the delusions of patients with delusional disorders. We highlight that developing evidence-based tests or benchmark sets to anticipate the behavior of the language models in collaboration with MHPs is critical when leveraging LLMs for clinical mental health settings.
Incorporating Perspectives of MHPs in Testing and Monitoring. Considering the caveats of current LLMs (Korngiebel and Mooney, 2021), it is critical to involve MHPs when deploying LLM-driven systems for patients in mental health contexts. While planning the field deployment study of MindfulDiary, we identified specific roles that MHPs could play. In the pre-use phase, MHPs should determine the suitability of users and facilitate the onboarding process with patients. During the mid-use phase, they should closely monitor interactions with the LLM and be prepared to intervene in cases of crises or unexpected use scenarios. Furthermore, they can offer or adjust treatments periodically based on long-term data. Additionally, they should regularly re-evaluate the continued use of the system. While some of these tasks should carefully be designed not to burden MHPs too much, it is important that LLMs do not make autonomous decisions about patients (e.g., diagnosis, prescription, or crisis management) but instead operate under professional oversight.
Providing Safeguards for Hallucinated LLM Generations. Our clinician dashboard provided various summarized information, such as word cloud, aggregating multiple dialogue entries so that the clinician quickly grasps the gist of the dialogues. Although we underwent intensive testing with the LLM-driven data summarizer, the LLM-driven data processing may still suffer from inaccuracies, biases, and misinterpretation (Rawte et al., 2023; Ji et al., 2023) of patient sentiments or context, which could adversely affect treatment decisions and patient well-being. To mitigate such drawbacks of LLMs in our study, we provided sufficient guidance to MHPs, cautioning them that the LLM-generated information they receive may be error-prone. However, in real-world settings, MHPs might accept the outputs of LLM without much attention. Therefore, when involving LLM-driven data processing, the system should foster careful reviewing of the content based on the expertise of MHPs. For example, future systems could incorporate features like highlighting in vivo phrases that were directly mentioned by patients and signify key aspects of their experience and feelings. By contrasting the in vivo phrases with the LLM’s original text, the system can encourage MHPs to put more scrutiny on the LLM’s original interpretation, which may contain errors, and the actual inputs spoken by patients.
@@ 7.4. Limitations and Future Work
Our recruitment method could impact the generalizability of our findings, as we recruited the patient participants for our field study from a single university hospital. Although we aimed to recruit patients with diverse types and levels of symptoms, our participants are not representative samples of psychiatric patients. They were young (mostly adolescents) and consulted by a fixed number of psychiatrists. While this work is just a first step toward designing an LLM-driven journaling app for psychiatric patients, further investigation is necessary with subjects from various backgrounds. To implement our pipeline, we used OpenAI’s GPT API, which provided the most capable LLM at the time of our study and was accessible via commercial API. As GPT models are continually updated, later models may not yield the same conversational behavior. To generalize the performance of our conversational pipeline design, future work is needed to compare multiple versions of MindfulDiary with different underlying LLMs.
@@ Acknowledgements.
We thank our study participants for their time and efforts. We also thank Eunkyung Jo and Yubin Choi for providing feedback on the early draft of this paper. This work was supported through a research internship at NAVER AI Lab, and in part supported by a grant of the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (grant number : HI22C1962) and by the National Research Foundation of Korea grant funded by the Ministry of Science and ICT (No. 2022R1G1A1009100).